OpenStudy (anonymous):

The table below shows two equations: Equation 1 |4x – 3|− 5 = 4 Equation 2 |2x + 3| + 8 = 3 Which statement is true about the solution to the two equations? Equation 1 and equation 2 have no solutions. Equation 1 has no solution and equation 2 has solutions x = -4, 1. The solutions to equation 1 are x = 3, -1.5 and equation 2 has no solution. The solutions to equation 1 are x = 3, -1.5 and equation 2 has solutions x = -4, 1.

OpenStudy (anonymous):

\[|4x – 3|− 5 = 4\] start by adding 5 to get \[|4x-3|=9\] then solve \[4x-3=9\\ 4x=12\\ x=3\] and \[4x-3=-9\\ 4x=-6\\ x=-\frac{6}{4}=-\frac{3}{2}\]

OpenStudy (anonymous):

\[|2x + 3| + 8 = 3\] start by subtracting \(8\) to get \[|2x+3|=-5\] which has no solution because the absolute value of anything is always greater than or equal to zero, and therefore cannot be \(-5\)

OpenStudy (anonymous):

i would go with this one The solutions to equation 1 are x = 3, -1.5 and equation 2 has no solution.

OpenStudy (anonymous):

thanks i appreciate it! can you help with a few more?

OpenStudy (anonymous):

@satellite73